The weights of 1,000 children, in average, is 61kg with standard deviation of 10kg. Suppose the weights are normally distributed, how many children weigh less than 60kg?


1
Expert's answer
2022-02-07T16:03:13-0500

Let "X=" weight: "X\sim N(\mu, \sigma^2 )."

Given "\mu=61 \ kg, \sigma=10\ kg."

"P(X<60)=P(Z<\dfrac{60-61}{10})"

"=P(Z<-0.1)\approx 0.460172"

460 children weigh less than 60kg.